OpenStudy (anonymous):

Light travels in a vacuum at a speed of roughly 2.998x10^8 m/s and its speed in air is only negligible (0.03%) slower than that. a)How long does it take light to traverse 1.0 ft? b) That means when you look at something 1000 m away, you are seeing it how many seconds back in time?

OpenStudy (kira_yamato):

(a) 1 feet = 0.3048m time taken = 0.3048 / (2.998 x 10^8) ≈ 1.02 x 10^–9 (b) ∆ time = 1000 / (2.998 x 10^8) ≈ 3.34 x 10^–6